OpenStudy (anonymous):

whats the quotient of 22/7 % 11/14

Directrix (directrix):

I don't understand what this says: 22/7 % 11/14. Do you want to take 22/7 percent of 11/14 or is the question something else? @grapehead

OpenStudy (anonymous):

% in somethings means modulus so if its like that then this would be 1/7 modulus 11/14, which is just (1/7)/(11/14)=14/77? (ithink)

Directrix (directrix):

Live and learn. I had not seen that before. Thanks for clarification. Somebody posted that the question translates to: 22/7 + 11/14 = whats the quotient? That looks like a "find the sum" problem to me.

OpenStudy (anonymous):

the answer choices arent the answer im seeing

Directrix (directrix):

@grapehead What are the answer choices?

OpenStudy (anonymous):

a.8 b.4 c.1/2 d.1/4

Directrix (directrix):

@Axohatah Could you help?

Directrix (directrix):

If 22/7 % 11/14 means 22/7 divided by 11/14, then 22/7 divided by 11/14 = 22/7 times 14/11 = ?

Directrix (directrix):